40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Apr 29 22:53:06 2018, gautam, Update, General, DARM actuation estimate darmEst.pdf
    Reply  Tue May 1 19:37:50 2018, gautam, Update, General, DARM actuation estimate darmEst.pdfdarmEst_time.pdf
       Reply  Mon May 7 16:23:06 2018, gautam, Update, General, DARM actuation estimate darmProj.pdf37.pngMCalign_20180507.png
          Reply  Mon May 7 20:01:14 2018, Rorpheus, Update, General, Use anti-dewhitening + show CARMA/DARMA 
             Reply  Thu May 10 14:13:22 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdf
                Reply  Fri May 11 13:58:42 2018, rana, Update, General, More refinement of DARM control signal projection 
                   Reply  Fri May 11 19:02:52 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdfDARMsuppression.pdfControlLoop.pdf
                      Reply  Sat May 12 10:02:03 2018, rana, Update, General, More refinement of DARM control signal projection 
                         Reply  Sun May 20 17:43:01 2018, rana, Update, Electronics, How to choose resistors 
                      Reply  Mon May 14 18:58:32 2018, Kevin, Update, PonderSqueeze, Squeezing with no SRM homodyne_heatmap.pdfdisplacement_noise.pdfnoise_budget.pdf
                         Reply  Tue Oct 2 23:57:16 2018, gautam, Update, PonderSqueeze, Squeezing scenarios PonderSqueeze_NB_LORIN.pdf
                   Reply  Tue May 15 21:56:57 2018, gautam, Update, General, Stack measurement setup decommissioned 
                      Reply  Thu May 17 09:14:38 2018, Steve, Update, General, Stack measurement setup decommissioned Load_Cell_Measurement_Set_Up.jpgETMX_stack_up_down.png
Message ID: 13822     Entry time: Mon May 7 16:23:06 2018     In reply to: 13805     Reply to this: 13823
Author: gautam 
Type: Update 
Category: General 
Subject: DARM actuation estimate 

Summary:

Using the Wiener Filter estimate of the DARM disturbance we will have to cancel, I computed how the control signal would look like for a few scenarios. Our DACs are 16-bit, +/-10V (i.e. +/-32,768cts-pk, or ~23,000cts RMS). We need to consider the shape of the de-whitening filter to conclude whether it is feasible to increase the series resistance by x10 or not.

Some details:

Note that in this first computation, I have not considered

  • Actuation range required by other loops (e.g. local damping, Oplev etc).
  • At some point, I need to add the 2P/c radiation pressure disturbance as well.
  • The control signal is calculated assuming we are actuating equally on both ETMs (but with opposite phase).
  • RMS computation is done from 30 Hz downwards, as above 30 Hz, I think the estimate from the previous elog is not true seismic displacement. 
  • De-whitening filters (or digital whitening), which will be required to suppress DAC noise at 100Hz.
  • DARM loop shape, specifically low-pass to avoid sensing noise injection. In this calculation, I just used the pendulum transfer function.

While doing this calculation, I have accounted for the fact that right now, the analog de-whitening filters in the ETM drive chain have a x3 gain which we will remove. Actually this is an assumption, I have not yet measured a transfer function, maybe I'll do one channel at EY to confirm. Also, the actuator gains themselves need to be confirmed.

As I was looking at the coil driver schematic more closely, I realized that there are actually two separate series resistances, one for the fast controls path, and another for the DC bias voltage from the slow ADCs. So I think we have been underestimating the Johnson noise of the coil drivers by sqrt(2). I've also attached screenshots of the IFOalign and MCalign screens. The two  ITMs and ETMX have pitch DC bias values that are compatible with a x10 increase of the series resistance. But even so, we will have ~3pA/rtHz per coil from the two resistances.


gautam 8pm May8: Seems like I had confirmed the x3 gain in the EX de-whitening board when Johannes and I were investigating the AI board offset.

Attachment 1: darmProj.pdf  99 kB  Uploaded Mon May 7 18:01:10 2018  | Hide | Hide all
darmProj.pdf
Attachment 2: 37.png  20 kB  Uploaded Mon May 7 18:02:58 2018  | Hide | Hide all
37.png
Attachment 3: MCalign_20180507.png  7 kB  Uploaded Mon May 7 18:03:03 2018  | Hide | Hide all
MCalign_20180507.png
ELOG V3.1.3-